OpenStudy (anonymous):

Based on the graph of y = x2 − 2x − 3, what is the x coordinate of the positive x-intercept

OpenStudy (anonymous):

Put y = 0 here.. and find x first..

OpenStudy (anonymous):

\[x^2 - 2x - 3 = 0\] Factorize this..

OpenStudy (anonymous):

(x + 1)(x - 3) x + 1 = 0 / x = -1 X - 3 =0 / x + 3 ???

OpenStudy (anonymous):

x =3

OpenStudy (anonymous):

So what is positive one?/ -1 or 3 ??

OpenStudy (anonymous):

3!

OpenStudy (anonymous):

So positive is 3 So this will be x intercept that is positive..

OpenStudy (anonymous):

And this will be the x coordinate of positive x intercept..

OpenStudy (anonymous):

ok thanks you :))

OpenStudy (anonymous):

Welcome dear..